Records indicate that , on average , 3.2 breakdowns per day occur on an urban highway during the morning rush hour. Assume that the distribution is poisson. Find the probability that on any given day there will be more than four breakdowns on this highway during the morning rush hour.
Answer by Boreal(15235)

probability of each breakdown from 0 to 3 is e^-lambda*lambda^x/x!
so p(0)=e^-3.2=0.0408
p(1)=).0408*3.2^1=0.1304 (round at end)
p(2)=0.0408*3.2^2/2!=0.2089
p(3)=0.0408*3.2^3/6=0.2228
the answer is 1-that sum=0.3971 (0.3975 with a calculator with rounding at the end). Find out where the intermediate steps should be rounded.